What is the value of "x" in the quadratic equation X squared -8x+12=0
Answer by amarjeeth123(569)

x^2-8x+12=0
x^2-6x-2x+12=0
x(x-6)-2(x-6)=0
(x-2)(x-6)=0
x=2 or x=6